# How AI Ideas Affect the Creativity, Diversity, and Evolution of Human Ideas: Evidence From a Large, Dynamic Experiment

### v2

@@ Abstract.
artificial intelligence, creativity, large language models, cultural evolution
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Preprint; Jul 03, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Human-centered computing Collaborative interaction
@@ 1. Introduction
If we think of culture as a “loop” where individuals and societies shape each other through exchanges of ideas and practices (Richerson and Boyd, 2008; Boyd and Richerson, 1988), then a question emerges: What happens when generative AI enters the “culture loop?” Exposure to LLMs (large language models) is increasing rapidly: When released, ChatGPT was the fastest-growing consumer application in history (Hu, 2023). Moreover, we are likely exposed to even more AI content than we realize: Humans overestimate their ability to distinguish AI from human content (Jakesch et al., 2022). This exposure likely matters: Ideas we see affect the ideas we create (Nijstad and Stroebe, 2006). How will the rapid rise of exposure to LLM-generated ideas affect the creativity, diversity, and evolution of human ideas? And to what extent do AI ideas influence human ideas?
The scale of ‘passive exposure’ to AI ideas is high, and different from prior human-AI interactions. By ‘passive exposure’, we refer to cases when (A) users see LLM outputs but do not have an active role in the creation of these outputs and (B) users are given no instructions to actively engage with these outputs. ‘Passive exposure’ approximates how users often encounter LLM outputs in the real world. For example, OpenAI users generate 100 billion words per day (Griffin, 2024). It is likely that the number of people who are merely seeing (i.e., passively exposed) AI output is significantly larger than the number of people who are creating (i.e., actively engaging) with these systems. Arguably, many future human-AI teams will exhibit this relationship. Yet in existing studies of human-AI creativity, participants are often actively interacting with an AI system (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019; Padmakumar and He, 2024). It is crucial, then, to understand how passive AI exposure shapes human ideas.
As AI exposure has increased, so have concerns over AI disclosure (Hancock et al., 2020) (whether providers should disclose when they use AI systems). California, for instance, considered passing a law requiring disclosure on behalf of anyone using bots on social media (Williams, 2018). Concerns regarding the disclosure of LLMs are only more likely to grow. In domains ranging from poetry (Köbis and Mossink, 2021) to online social media profiles (Jakesch et al., 2022), LLM output is increasingly indistinguishable from that of humans. We are interested, then, if disclosing ideas as coming from AI moderates the effect of AI exposure.
@@
Motivated by these dynamics, we conducted a large-scale experiment to systematically test how AI exposure and disclosure affect the creativity, diversity, and evolution of human ideas. We employ a variant of the Alternate Uses Task (AUT, (Guilford, 1978)), a common measure of creativity, and manipulate exposure to LLM ideas. In the AUT, participants are told to think of non-obvious uses of an item. For example: What is a creative use for a tire? In our variant, participants complete the AUT for an item after viewing example ideas. These examples constitute our manipulation. Examples vary in AI exposure (none, few, or many AI examples) and AI disclosure (whether AI-generated ideas are labeled as such). The human-generated ideas in each example set come from prior participants in the same experimental condition. See Figure 1 for a graphical depiction of the experiment.
Our dynamic experiment design—ideas from prior participants are used as stimuli for future participants—speaks to the interdependent process of cultural creation: creative ideas are built upon prior ideas. Hence, we capture the compounding effects of having LLMs “in the culture loop”. It is also intended to mimic possible futures for human-AI teams. Our design allows us to observe not just average levels but also temporal dynamics of creativity and diversity in each condition. Taken together, our results provide insights into the role of LLMs in shaping collective thought.
Concretely, our main findings are:
@@
    High AI exposure increases collective diversity but not individual creativity. We find that high AI exposure increases collective idea diversity, but does not affect individual creativity. Our high-powered null finding around creativity can inform public debates over the creative impact of seeing AI ideas ‘in the wild’. However, we found conditions with high levels of AI exposure had more collective idea diversity. That is, ideas in the high AI exposure conditions were more different from each other. Our findings around creativity and diversity suggest the effect of AI exposure may be nuanced: The introduction of AI ideas into human society may yield more diverse but no better human ideas.
@@
    High AI exposure increases the speed at which idea diversity develops. Culture is constantly evolving (Boyd et al., 2011; Boyd and Richerson, 1988), yet many laboratory experiments are not designed to model this evolution. Through our dynamic design, we find that high AI exposure increases not only the average levels of collective idea diversity but also the rate of change in idea diversity. This is a consequential finding since even small differences in rates of change can lead to large cumulative differences over time.
@@
    People who identify as creative are less influenced by AI disclosure. Prior work argues that attitudes and expectations shape engagement with human-AI co-creation systems (Gero et al., 2022). Due to our large sample size, we can model this heterogeneity. We find that for users who self-identify as highly creative, adoption of AI ideas is not influenced by AI disclosure. But AI disclosure did affect the adoption of AI ideas for users who self-identified as low in creativity. This finding suggests that highly creative people will not be “duped” into adopting AI ideas.
@@
    Participants may adopt AI ideas for harder prompts. We find that when AI ideas are disclosed, participants are more likely to adopt the ideas of AI for difficult AUT prompts. This suggests that users will rely on AI ideas not for trivial creative tasks but for difficult ones. But since this finding is based on a small number of prompts, we view this finding as speculative and encourage more work on the topic.
@@ 1.1.1. Creativity
Creativity is defined in many ways (Walia, 2019). But one common conception is divergent thinking (Guilford, 1967). This is when “an individual solves a problem or reaches a decision using strategies that deviate from commonly used or previously taught strategies” (of Psychology, [n. d.]). One of the most common (Abraham, 2016) tests of divergent thinking is the Alternate Uses Task (AUT) (Guilford, 1978) [^1] https://www.mindgarden.com/67-alternate-uses, where participants are asked to think of an original use for an everyday object. Traditionally, responses to the AUT are measured along four dimensions: originality (how original the idea is), elaboration (how much the participant elaborates on the idea), fluency (how many ideas), and flexibility (different categories of ideas). The latter two can only be measured if the participant provides multiple responses to the same question. Due to our research design [^2] Participants see the most recent responses in the condition as stimuli, so if one participant brainstorms many responses that participant would be over-represented in future participants’ example sets., we have participants generate just one creative idea (as in (Beaty et al., 2022)), and we focus on originality.
We follow a long tradition of scoring responses to the AUT computationally (Yu et al., 2023; Beaty and Johnson, 2021; Beaty et al., 2022; Yang et al., 2023; Organisciak et al., 2022; Dumas et al., 2021). Specifically, we measure the creativity of AUT ideas with an existing fine-tuned GPT-3 classifier (Organisciak et al., 2022, 2023), which has an r=0.81 overall correlation with human judgments of AUT originality. Moreover, we chose AUT items for our experiment where the classifier had the highest accuracy333tire (r=0.91), pants (r=0.91), shoe (r=0.91), table (r=0.9), and bottle (r=0.88). Note that our task is highly ‘in-domain’ for the classifier: we ask participants to do the same exact task for the same exact items the model was trained on. We refer to the originality score from this classifier as individual-level creativity, though we note that future work can explore other dimensions of creativity (such as fluency). We discuss this classifier in more detail in Section 3.
@@ 1.1.2. Idea Diversity & AI Adoption
In addition to creativity, we measure how our experimental factors (LLM exposure and LLM disclosure) shape the diversity of ideas that participants produce. This is a complementary measure to creativity. Creativity is often thought of as an individual-level outcome. Diversity is a collective outcome. Put another way, creativity is a property of an idea while diversity is a property of an idea set. We measure two sides of diversity—semantic divergence (which we refer to as idea diversity) and semantic convergence towards AI ideas (which we refer to as AI adoption).
@@
  • AI adoption is the maximum cosine similarity between the embedding of the idea a participant submits and the embeddings of AI examples that the participants see. Following Roemmele (2021), we use the max rather than a measure of central tendency because if a participant is inspired by an idea, it would likely be a single idea. As robustness checks, we also measure the mean and median pairwise similarity between the submitted idea and an AI example, but these are noisier measures of adoption.
@@ 2. Related Work
Our work bridges three research streams: human-AI co-creation, crowd-sourced creativity, and complex systems. AI ideas are scattered amongst human ideas, whether or not we can tell (Jakesch et al., 2022). This exposure presumably affects the ideas we create (co-creation). And our ideas presumably affect the ideas others create (crowdsourced creativity). While real-world culture is dynamic and evolving, most experiments are not set up to capture evolution (collective dynamics). By employing a large-N sample size and ‘many-worlds’ setup, we model the complex dynamics of AI influence. After discussing how our study bridges these streams, we turn to the particular kind of creativity and diversity our experiment captures and what is known about how our two factors (LLM exposure; LLM disclosure) would affect these outcome variables. However, much of the relevant literature gives conflicting predictions, a key motivation for conducting the current study.
2.1. Situating Our Work Between Co-Creation, Crowd Creativity, and Collective Dynamics
2.1.1. Human-AI Co-Creation
@@
As the creative ability of AI advances (Miller, 2019), researchers explored how co-creating with AI affects human creativity. Much of this research explores creative writing with language models, in particular (Gero et al., 2022; Mirowski et al., 2023; Lee et al., 2022; Yang et al., 2022; Yuan et al., 2022; Roemmele, 2021; Di Fede et al., 2022; Hitsuwari et al., 2022; Gero, 2023; Mizrahi et al., 2020; Gero and Chilton, 2019; Padmakumar and He, 2024). While most prior work in this domain involves users actively engaging with custom systems, our study is concerned with passive exposure to outputs from off-the-shelf models. (By ‘passive exposure’ we mean that (1) users are shown LLM outputs but did not have an active role in the creation of these outputs and that (2) users were given no instructions to actively engage with these outputs; they were merely shown the LLM outputs. ‘Passive exposure’ approximates how users often encounter LLM outputs in the real world.) Human-AI co-creation shows that the relationship between AI ideas and their effect on human creativity is nuanced, with task-level and attitudinal factors playing a role. Roemmele (2021) found seeing AI examples influenced outcomes for hard, but not easy, prompts. Gero et al. (2022) found the quality of LLM outputs did not correlate with perceived usefulness. This is consistent with other research showing large variance in the perceived usefulness of outputs from co-creation systems (Calderwood et al., 2020), suggesting human attitudes partially determine the utility of AI creativity aids. We extend this predominantly qualitative work with a large-scale quantitative study.
@@
Most similar to our work, several studies have explored how (post-ChatGPT) generative artificial intelligence affects creativity and diversity. Several studies found that ideating with generative AI can decrease diversity (Anderson et al., 2024; Doshi and Hauser, 2024; Padmakumar and He, 2024). Some studies suggest generative AI increases individual creativity (Doshi and Hauser, 2024; Dell’Acqua et al., 2023) while others (Anderson et al., 2024) find no effect. Our study offers several additions to this literature. First, because of its dynamic design, we test the long-run effects of AI, where ideas feed forward to future participants. Second, our study is concerned with “passive exposure”: Participants are not told ideas are from AI, and are not instructed to engage with these ideas. By systematically ablating whether AI ideas are disclosed as such, we can explore if the effect of AI ideas depends on knowledge of where the idea is from. Third, we employ a large sample size—which is useful since it provides power for precise estimates of effects and the ability to capture heterogeneity. Moreover, our large sample is comprised of creative professionals and technology-oriented users, two groups most relevant to the phenomenon.
@@ 2.1.2. Crowdsourced Creativity
Crowdsourcing can enhance creative outcomes (Yu and Nickerson, 2011, 2013; Nickerson and Sakamoto, 2010; Huang et al., 2020; Siangliulue et al., 2015). For example, Yu and Nickerson (2013) devised a method where crowds build on each other’s ideas by combining ideas from previous generations. Later generations of ideas were rated as more creative compared to earlier generations. Siangliulue et al. (2015) found that the creativity and diversity of idea sets that participants saw influenced the creativity and diversity of what these participants produced. This supports a main contention of our paper: AI exposure matters because the ideas we see affect the ideas we create. We incorporate elements of crowdsourced creativity, particularly in measuring how creativity and diversity unfold over subsequent generations.
2.1.3. Collective Dynamics & Many-Worlds Experiments
Prior work in complex systems and computational sociology highlights the importance of studying collectives to understand social dynamics (Salganik and Watts, 2009). Meanwhile, traditional experiments focus on individuals. Identifying the effect of AI ideas on the diversity and evolution human ideas similarly requires an examination of complex systems as opposed to individuals in isolation. Hence, our experimental design draws on the “many-world” paradigm (e.g., (Salganik et al., 2006)):We create multiple, parallel realizations of worlds with and without AI ideas, each evolving independently under controlled conditions. By employing a large-N sample size and many different parallel worlds, we can better understand the collective effects of AI ideas on human ideas. Beyond simple averages, we can also model how AI ideas affect the evolution of human ideas.

### v3

@@ Abstract.
artificial intelligence, large language models, human-computer interaction, creativity, collective intelligence, cultural evolution
journalyear: 2025copyright: ccconference: Collective Intelligence Conference; August 4–6, 2025; San Diego, CA, USAbooktitle: Collective Intelligence Conference (CI 2025), August 4–6, 2025, San Diego, CA, USAdoi: 10.1145/3715928.3737481isbn: 979-8-4007-1489-4/2025/08ccs: Human-centered computingccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Empirical studies in collaborative and social computingccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Human-centered computing Collaborative interaction
@@ 1. Introduction
If we think of culture as a feedback loop where individuals and societies shape each other through exchanges of ideas and practices (Richerson and Boyd, 2008; Boyd and Richerson, 1988), then what happens when generative AI enters this “culture loop”? Exposure to LLMs (large language models) is increasing: When released, ChatGPT was the fastest-growing consumer application in history (Hu, 2023). Moreover, we are likely exposed to even more AI content than we realize: Humans overestimate their ability to distinguish AI from human content (Jakesch et al., 2022). This exposure likely matters: Ideas we see affect the ideas we create (Nijstad and Stroebe, 2006). How, then, will the rapid rise of exposure to LLM-generated ideas affect the creativity, diversity, and evolution of human ideas? And to what extent do AI ideas influence human ideas? These questions have implications for how AI may shape collective intelligence (Peeters et al., 2021; Cui and Yasseri, 2024; Burton et al., 2024). Our study speaks to these questions in the specific context of brainstorming, an activity that occurs across many domains.
The widespread dissemination of AI-generated content creates a phenomenon we call “passive exposure,” which differs from the “active engagement” primarily studied in prior human-AI interactions. By “passive exposure”, we refer to cases when (A) users see LLM outputs but do not have an active role in the creation of these outputs and (B) users are given no instructions to actively engage with these outputs. Passive exposure approximates how users often encounter LLM outputs in the real world. For example, OpenAI users generate 100 billion words per day (Griffin, 2024). It is likely that the number of people who are merely seeing (i.e., passively exposed) AI output is significantly larger than the number of people who are creating (i.e., actively engaging) with these systems. Yet in existing studies of human-AI creativity, participants are often actively interacting with an AI system (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019; Padmakumar and He, 2024).
As AI exposure has increased, so have concerns over AI disclosure (Hancock et al., 2020) (whether providers should disclose when they use AI systems). California considered requiring disclosure on behalf of anyone using bots on social media (Williams, 2018). Concerns regarding the disclosure of LLMs are only more likely to grow. LLM output is increasingly indistinguishable (Jakesch et al., 2022) from that of humans. We are interested if disclosing ideas as coming from AI moderates the effect of AI exposure.
@@
Motivated by these dynamics, we conducted a large-scale experiment to systematically examine how AI exposure and disclosure influence the creativity, diversity, and evolution of human ideas. We utilize a variant of the Alternate Uses Task (AUT, (Guilford, 1978)), a standard measure of creativity, and manipulate exposure to LLM-generated ideas. In the AUT, participants generate non-obvious uses for a common item. For example: What is a creative use for a tire? In our variant, participants complete the AUT after viewing example ideas, which serve as our experimental manipulation. Examples vary in AI exposure (none, few, or many AI examples) and AI disclosure (whether AI-generated ideas are explicitly labeled). The human-generated ideas in each example set originate from prior participants in the same experimental condition. See Figure 1 for a depiction.
Our dynamic experiment design leverages ideas from prior participants as stimuli for future participants, reflecting the interdependent nature of brainstorming. Ideas build upon previous ideas. This approach captures (certain aspects of) the compounding effects of integrating LLMs into the “culture loop” and simulates potential futures for human-AI collaboration. Our design enables us to observe both average performance levels and temporal dynamics of creativity and diversity across conditions. This setup provides insights into how LLMs shape collective thought.
1.1. Findings
Our experiment manipulates two key factors: (1) AI exposure (the number of AI-generated ideas in example sets) and (2) AI disclosure (whether AI ideas are labeled as such). We investigate three outcomes: individual creativity, idea diversity (measured at individual and collective levels), and AI idea adoption (who and when adopts AI example ideas). For creativity and diversity, we analyze both average levels and temporal evolution across experimental iterations. See Table 1 for our measures. Here are our findings.
@@ 1.1. Findings
    High AI exposure increases collective diversity but not individual creativity. We find that AI exposure did not affect individual creativity. However, conditions with high levels of AI exposure had more collective idea diversity. That is, ideas in the high AI exposure conditions were more different from each other but not necessarily better. Our findings around creativity and diversity suggest the effect of AI exposure may be nuanced.
@@
    High AI exposure increases the speed at which idea diversity develops. Through our dynamic design, we find that high AI exposure increases not only the average levels of collective idea diversity but also the rate of change in idea diversity. This is a consequential finding since even small differences in rates of change can lead to large cumulative differences over time.
@@
    People who identify as creative are less influenced by AI disclosure. We find that for users who self-identify as highly creative, adoption of AI ideas is not influenced by AI disclosure. But AI disclosure did affect the adoption of AI ideas for users who self-identified as low in creativity. This suggests that highly creative people will not be “duped” into adopting AI ideas.
@@
    Participants may adopt AI ideas for harder prompts. We find that when AI ideas are disclosed as such, participants were more likely to adopt the ideas of AI for difficult AUT prompts. This suggests that users may rely on AI ideas not for trivial creative tasks but for difficult ones. Since this finding is based on a small number of AUT items, this is speculative.
@@ 1.2.1. Creativity
Creativity is defined in many ways (Walia, 2019). But one common conception is divergent thinking (Guilford, 1967): When “an individual solves a problem or reaches a decision using strategies that deviate from commonly used or previously taught strategies” (of Psychology, [n. d.]). One of the most common (Abraham, 2016) tests of divergent thinking is the Alternate Uses Task (AUT) (Guilford, 1978) [^1] https://www.mindgarden.com/67-alternate-uses, where participants are asked to think of an original use for an everyday object. Traditionally, responses to the AUT are measured along four dimensions: originality (how original the idea is), elaboration (how much the participant elaborates on the idea), fluency (how many ideas), and flexibility (different categories of ideas). The latter two can only be measured if the participant provides multiple responses to the same question. Due to our research design [^2] Participants see the most recent responses in the condition as stimuli, so if one participant brainstorms many responses, that participant would be over-represented in future participants’ example sets., we have participants generate just one creative idea (as in (Beaty et al., 2022)), and we focus on originality.
We follow a tradition of scoring responses to the AUT computationally (Yu et al., 2023; Beaty and Johnson, 2021; Beaty et al., 2022; Yang et al., 2023; Organisciak et al., 2023; Dumas et al., 2021). Specifically, we measure the creativity of AUT ideas with an existing fine-tuned GPT-3 classifier (Organisciak et al., 2022), which has an r=0.81 overall correlation with human judgments of AUT originality. Moreover, we chose AUT items for our experiment where the classifier had the highest accuracy333tire (r=0.91), pants (r=0.91), shoe (r=0.91), table (r=0.9), and bottle (r=0.88). Our task is highly ‘in-domain’ for the classifier: we ask participants to do the same exact task for the same exact items the model was trained on. We refer to the originality score from this classifier as individual-level creativity, but future work can explore other dimensions of creativity (such as fluency). We discuss this classifier more in Appendix A.1.
@@ 1.2.2. Idea Diversity & AI Adoption
In addition to creativity, we measure how our experimental factors (LLM exposure and LLM disclosure) shape the diversity of ideas that participants produce. This is a complementary measure to creativity. Creativity is often thought of as an individual outcome. Diversity is a collective outcome. Equivalently, creativity is a property of an idea while diversity is a property of an idea set. We measure two sides of diversity—semantic divergence (which we refer to as idea diversity) and semantic convergence towards AI ideas (which we refer to as AI adoption).
@@
  • AI adoption is the maximum cosine similarity between the embedding of the idea a participant submits and the embeddings of AI examples that the participants see. Following (Roemmele, 2021), we use the max rather than a measure of central tendency because if a participant is inspired by an idea, it would likely be a single idea. As robustness checks, we also measure the mean and median pairwise similarity between the submitted idea and an AI example, but these are noisier measures of adoption.
@@ 2. Related Work
Our work bridges three research streams: human-AI co-creation, crowd-sourced creativity, and collective dynamics. AI ideas are scattered amongst human ideas, whether or not we can tell (Jakesch et al., 2022). This exposure presumably affects the ideas we create (co-creation). And our ideas presumably affect the ideas others create (crowdsourced creativity). While real-world culture is dynamic and evolving, most experiments are not set up to capture evolution (collective dynamics). By employing a large-N sample size and “multiple-worlds” setup, we model the complex dynamics of AI influence. In Appendix C, we provide a more detailed overview of how our experimental factors (LLM exposure and disclosure) may affect creativity and diversity of ideas, as well as how attitudinal variables may moderate these effects. Overall, predictions from prior work and theory are conflicting—motivating the present study.
2.1. Human-AI Co-Creation
@@
As AI becomes more creative (Miller, 2019), researchers have explored how co-creating with AI affects human creativity. Much of this research explores creative writing with language models (Gero et al., 2022; Mirowski et al., 2023; Lee et al., 2022; Yang et al., 2022; Yuan et al., 2022; Roemmele, 2021; Di Fede et al., 2022; Hitsuwari et al., 2022; Gero, 2023; Mizrahi et al., 2020; Gero and Chilton, 2019; Padmakumar and He, 2024). While most prior work involves users actively engaging with custom systems, our study is concerned with passive exposure to off-the-shelf models. The relationship between AI ideas and their effect on human creativity is nuanced. Task-level and attitudinal factors play a role. For example, seeing AI examples influenced outcomes for hard, but not easy, prompts in Roemmele (2021), and Gero et al. (2022) found the quality of LLM outputs did not correlate with perceived usefulness. Large variance in the perceived usefulness of outputs from co-creation systems (Calderwood et al., 2020) suggests human attitudes partially determine the utility of AI creativity aids. We extend this predominantly qualitative work with a large-scale quantitative study.
@@
Several studies, like ours, explored how (post-ChatGPT) generative artificial intelligence affects creativity and diversity. Some studies show that ideating with generative AI decreases diversity (Anderson et al., 2024; Doshi and Hauser, 2024; Padmakumar and He, 2024). Other studies suggest generative AI increases individual creativity (Doshi and Hauser, 2024; Dell’Acqua et al., 2023) while others find mixed effects on creativity (Anderson et al., 2024). Our study offers several additions. First, because of its dynamic design, we test the long-run effects of AI, where ideas feed forward to future participants. Second, our study is concerned with “passive exposure”: Participants are not told ideas are from AI, and are not instructed to engage with these ideas. By systematically ablating whether AI ideas are disclosed as such, we can explore if the effect of AI ideas depends on knowledge of where the idea is from. Third, we employ a large sample size—which is useful since it provides power for precise estimates of effects and the ability to capture heterogeneity. Fourth, our large sample is comprised of creative professionals and technology-oriented users, two groups relevant to the phenomenon.
@@ 2.1.1. Crowdsourced Creativity
Crowdsourcing can enhance creative outcomes (Yu and Nickerson, 2011, 2013; Nickerson and Sakamoto, 2010; Huang et al., 2020; Siangliulue et al., 2015). Siangliulue et al. (2015) found that the creativity and diversity of idea sets that participants saw influenced the creativity and diversity of what these participants produced. This supports a main contention of our paper. AI exposure matters because the ideas we see affect the ideas we create. We incorporate elements of crowdsourced creativity, particularly in measuring how creativity and diversity unfold over subsequent generations.
2.1.2. Collective Dynamics & Multiple-Worlds Experiments
Prior work in complex systems and sociology highlights the importance of studying collectives to understand social dynamics (Salganik and Watts, 2009). Meanwhile, traditional experiments focus on individuals. Identifying the effect of AI ideas on the diversity and evolution human ideas similarly requires an examination of complex systems as opposed to individuals in isolation. Hence, our experimental design draws on the “multiple-world” paradigm (e.g., (Salganik et al., 2006; Salganik and Watts, 2009)): We create multiple, parallel realizations of worlds with and without AI ideas, each evolving independently under controlled conditions. By employing a large-N sample size and many different parallel worlds, we can better understand the collective effects of AI ideas on human ideas. Beyond simple averages, we can also model how AI ideas affect the evolution of human ideas.